# QECO: A QoE-Oriented Computation Offloading Algorithm based on Deep Reinforcement Learning for Mobile Edge Computing

### v2

@@
QECO: A QoE-Oriented Computation Offloading Algorithm based on Deep Reinforcement Learning for Mobile Edge Computing
[…]
@@
In the realm of mobile edge computing (MEC), efficient computation task offloading plays a pivotal role in ensuring a seamless quality of experience (QoE) for users. Maintaining a high QoE is paramount in today’s interconnected world, where users demand reliable services. This challenge stands as one of the most primary key factors contributing to handling dynamic and uncertain mobile environment. In this study, we delve into computation offloading in MEC systems, where strict task processing deadlines and energy constraints can adversely affect the system performance. We formulate the computation task offloading problem as a Markov decision process (MDP) to maximize the long-term QoE of each user individually. We propose a distributed QoE-oriented computation offloading (QECO) algorithm based on deep reinforcement learning (DRL) that empowers mobile devices to make their offloading decisions without requiring knowledge of decisions made by other devices. Through numerical studies, we evaluate the performance of QECO. Simulation results validate that QECO efficiently exploits the computational resources of edge nodes. Consequently, it can complete 14% more tasks and reduce task delay and energy consumption by 9% and 6%, respectively. These together contribute to a significant improvement of at least 37% in average QoE compared to an existing algorithm.
@@ I Introduction
M  OBILE edge computing (MEC) [1] has emerged as a promising technological solution to overcome the challenges faced by mobile devices (MDs) when performing high computational tasks, such as real-time data processing and artificial intelligence applications [2] [3]. In spite of the MDs’ technological advancements, their limited computing power and battery may lead to task drops, processing delays, and an overall poor user experience. By offloading intensive tasks to nearby edge nodes (ENs), MEC effectively empowers computation capability and reduces the delay and energy consumption. This improvement enhances the users’ QoE, especially for time-sensitive computation tasks [4] [5].
Efficient task offloading in MEC is a complex optimization challenge due to the dynamic nature of the network and the variety of MDs and servers involved [6] [7]. In particular, determining the optimal offloading strategy, scheduling the tasks, and selecting the most suitable EN for task offloading are the main challenges that demand careful consideration. Furthermore, the uncertain requirements and sensitive latency properties of computation tasks pose nontrivial challenges that can significantly impact the computation offloading performance in MEC systems with limited resources.
I-A Related Work
To cope with the dynamic nature of the network, recent research has proposed several task offloading algorithms using machine learning methods. In particular, deep reinforcement learning (DRL) hold promises to determine optimal decision-making policies by capturing the dynamics of environments and learning strategies for accomplishing long-term objectives [8]. DRL can effectively tackle the challenges of MEC arising from the ever-changing nature of networks, MDs, and servers’ heterogeneity. This ultimately improves the MD users’ QoE. In [9], Huang et al. focused on a wireless-powered MEC. They proposed a DRL-based approach, capable of attaining near-optimal decisions. This is achieved by selectively considering a compact subset of candidate actions in each iteration. In [10], the authors proposed an offloading algorithm using deep Q-learning for wireless-powered Internet of Things (IoT) devices in MEC systems. This algorithm aims to minimize the task drop rate while the devices solely rely on harvested energy for operation. In [11], Zhao et al. proposed a computation offloading algorithm based on DRL, which addresses the competition for wireless channels to optimize long-term downlink utility. In this approach, each MD requires quality-of-service information from other MDs. Tang et al. in [12] investigated the task offloading problem for indivisible and deadline-constrained computational tasks in MEC systems. The authors proposed a distributed DRL-based offloading algorithm designed to handle uncertain workload dynamics at the ENs. Sun et al. in [13] explored both computation offloading and service caching problems in MEC. They formulated an optimization problem that aims to minimize the long-term average service delay. They then proposed a hierarchical DRL framework, which effectively handles both problems under heterogeneous resources. Dai et al. in [14] introduced the integration of action refinement into DRL and designed an algorithm to concurrently optimize resource allocation and computation offloading. In [15], Huang et al. proposed a DRL-based method based on a partially observable MDP, which guarantees the deadlines of real-time tasks while minimizing the total energy consumption of MDs. Liu et al. in [16] investigated a two-timescale computing offloading and resource allocation problem and proposed a resource coordination algorithm based on multi-agent DRL, which can generate interactive information along with resource decisions. Zhou et al. in [17] used an MDP to study MEC and modeled the interactions of the environment. They proposed a Q-learning approach to achieve optimal resource allocation strategies and computation offloading. In [18], Gao et al. introduced an attention-based multi-agent algorithm designed for decentralized computation offloading. This algorithm effectively tackles the challenges of dynamic resource allocation in large-scale heterogeneous networks. Gong et al. in [19] proposed a DRL-based network structure in the industrial IoT systems to jointly optimize task offloading and resource allocation in order to achieve lower energy consumption and decreased task delay. Liao et al. in [20] introduced a double reinforcement learning algorithm for performing online computation offloading in MEC. This algorithm optimizes transmission power and scheduling of CPU frequency when minimizing both task computation delay and energy consumption.
I-B Motivation and Contributions
Although DRL-based methods have demonstrated their effectiveness in handling network dynamics, task offloading still encounters several challenges that require further attention. QoE is a time-varying performance measure that reflects user satisfaction and is not affected only by delay, as assumed in [9][13], but also by energy consumption. Albeit some existing works such as [14][20], have investigated the trade-off between delay and energy consumption, they fail to properly address the user demands and fulfill QoE requirements. A more comprehensive approach is required to address the dynamic requirements of individual users in real-time scenarios with multiple MDs and ENs. In contrast to the aforementioned works [9][20], we propose a DRL-based distributed algorithm that provides users with an appropriate balance among QoE factors based on their demands. We also explore a more realistic MEC scenario involving delay-sensitive tasks with processing deadlines, posing a more intricate challenge.
In this study, we delve into the computation task offloading problem in MEC systems, where strict task processing deadlines and energy constraints can adversely affect the system performance. We propose a distributed QoE-oriented computation offloading (QECO) algorithm that leverages DRL to efficiently handle task offloading in uncertain loads at ENs. This algorithm empowers MDs to make offloading decisions utilizing only locally observed information, such as task size, queue details, battery status, and historical workloads at the ENs. By adopting the appropriate policy based on each MD’s specific requirements at any given time, the QECO algorithm significantly improves the QoE for individual users.
Our main contributions are summarized as follows:
  • Task Offloading Problem in the MEC System: We formulate the task offloading problem as an MDP for time-sensitive tasks. This approach takes into account the dynamic nature of workloads at the ENs and concentrates on providing high performance in the MEC system while maximizing the long-term QoE.
  • DRL-based Offloading Algorithm: To address the problem of long-term QoE maximization, we focus on task completion, task delay, and energy consumption to quantify the MDs’ QoE. We propose QECO algorithm based on DRL that empowers each MD to make offloading decisions independently, without prior knowledge of the other MDs’ tasks and offloading models. With a focus on the MD’s battery level, our approach leverages deep Q-network (DQN) [21] and long short-term memory (LSTM) [22] to prioritize and strike an appropriate balance between QoE factors. We also analyze the training convergence and complexity of the proposed algorithm.
  • Performance Evaluation: We conduct comprehensive experiments to evaluate the QECO’s performance as well as its training convergence under different computation workloads. The results demonstrate that our algorithm quickly converges and effectively utilizes the processing capabilities of MDs and ENs, resulting in substantial improvement of at least 37% in average QoE. This advantage is achieved through a 14% increase in the number of completed tasks, along with 9% and 6% reductions in task delay and energy consumption, respectively, when compared to the potential game-based offloading algorithm (PGOA) [23] and several benchmark methods.
The structure of this paper is as follows. Section II presents the system model, followed by the problem formulation in Section  III. In Section IV, we present the algorithm, while Section V provides an evaluation of its performance. Finally, we conclude in Section VI.
II System Model
We investigate a MEC system consisting of a set of MDs denoted by =$\mathcal{I}=\{1,2,...,I\}$, along with a set of ENs denoted by $\mathcal{J}=\{1,2,...,J\}$, where $I$ and $J$ represent the number of MDs and ENs, respectively. We regard time as a specific episode containing a series of $T$ time slots denoted by $\mathcal{T}=\{1,2,\ldots,T\}$, each representing a duration of $\tau$ seconds. As shown in Fig. 1, we consider two separate queues for each MD to organize tasks for local processing or dispatching to ENs, operating in a first-in-first-out (FIFO) manner. The MD’s scheduler is responsible for assigning newly arrived tasks to each of the queues at the beginning of the time slot. On the other hand, we assume that each EN $j\in\mathcal{J}$ consists of $I$ FIFO queues, where each queue corresponds to an MD $i\in\mathcal{I}$. When each task arrives at an EN, it is enqueued in the corresponding MD’s queue.
We define $z_{i}(t)$ as the index assigned to the computation task arriving at MD $i\in\mathcal{I}$ in time slot $t\in\mathcal{T}$. Let $\lambda_{i}(t)$ denote the size of this task in bits. The size of task $z_{i}(t)$ is selected from a discrete set $\Lambda=\{\lambda_{1},\lambda_{2},\ldots,\lambda_{\theta}\}$, where $\theta$ represents the number of these values. Hence, $\lambda_{i}(t)\in\Lambda\cup\{0\}$ to consider the case that no task has arrived. We also denote the task’s processing density as $\rho_{i}(t)$ that indicates the number of CPU cycles required to complete the execution of a unit of the task. Furthermore, we denote the deadline of this task by $\Delta_{i}(t)$ which is the number of time slots that the task must be completed to avoid being dropped.
II-A Communication Model
$\delta_{i}^{\text{T}}(t)={\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}i}\left[{\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}i}\max\limits_{t^{{}^{\prime}}{\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}i}\in{\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}i}\{0,1,\ldots,t-1\}}l_{i}^{\text{T}}{\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}i}(t^{{}^{\prime}})-t+1{\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}i}\right]^{+}{\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}i},$ (1)
where $[\cdot]^{+}=$ max$(0,\cdot)$ and $l_{i}^{\text{T}}(0)=0$ for the simplicity of presentation. Note that the value of $\delta_{i}^{\text{T}}(t)$ only depends on $l_{i}^{\text{T}}(t)$ for $t^{\prime}<t$. If MD $i\in\mathcal{I}$ schedules task $z_{i}(t)$ for dispatching in time slot $t\in\mathcal{T}$, then it will either be dispatched or dropped in time slot $l_{i}^{\text{T}}(t)$, which is
@@ II-A Communication Model
$D_{i}^{\text{T}}(t)=\sum_{\mathcal{J}}y_{i,j}(t){\lambda_{i}(t)\over r_{i,j}(t)\tau}.$ (3)

### v3

@@
QECO: A QoE-Oriented Computation Offloading Algorithm based on Deep Reinforcement
Learning for Mobile Edge Computing
Iman Rahmaty ID , Hamed Shah-Mansouri ID , Member, IEEE, and Ali Movaghar ID , Life Senior Member, IEEE Received 27 March 2024; revised 28 November 2024; accepted 25 March 2025. Date of publication 1 April 2025; date of current version 27 June 2025. Recommended for acceptance by Dr. Hongbo Jiang. (Corresponding author: Hamed Shah-Mansouri.)Iman Rahmaty and Ali Movaghar are with the Department of Computer Engineering, Sharif University of Technology, Tehran 14588-89694, Iran (e-mail: iman.rahmati@sharif.edu; movaghar@sharif.edu).Hamed Shah-Mansouri is with the Department of Electrical Engineering, Sharif University of Technology, Tehran 14588-89694, Iran (e-mail: hamedsh@sharif.edu).Digital Object Identifier 10.1109/TNSE.2025.3556809Source code available at https://github.com/ImanRHT/QECO
@@
In the realm of mobile edge computing (MEC), efficient computation task offloading plays a pivotal role in ensuring a seamless quality of experience (QoE) for users. Maintaining a high QoE is paramount in today’s interconnected world, where users demand reliable services. This challenge stands as one of the most primary key factors contributing to handling dynamic and uncertain mobile environments. In this study, we delve into computation offloading in MEC systems, where strict task processing deadlines and energy constraints can adversely affect the system performance. We formulate the computation task offloading problem as a Markov decision process (MDP) to maximize the long-term QoE of each user individually. We propose a distributed QoE-oriented computation offloading (QECO) algorithm based on deep reinforcement learning (DRL) that empowers mobile devices to make their offloading decisions without requiring knowledge of decisions made by other devices. Through numerical studies, we evaluate the performance of QECO. Simulation results reveal that compared to the state-of-the-art existing works, QECO increases the number of completed tasks by up to 14.4%, while simultaneously reducing task delay and energy consumption by 9.2% and 6.3%, respectively. Together, these improvements result in a significant average QoE enhancement of 37.1%. This substantial improvement is achieved by accurately accounting for user dynamics and edge server workloads when making intelligent offloading decisions. This highlights QECO’s effectiveness in enhancing users’ experience in MEC systems.
@@ I Introduction
M  OBILE edge computing (MEC) [1] has emerged as a promising technological solution to overcome the challenges faced by mobile devices (MDs) when performing high computational tasks, such as real-time data processing and artificial intelligence applications [2]. In spite of the MDs’ technological advancements, their limited computing power and energy may lead to task drops, processing delays, and an overall poor user experience. By offloading intensive tasks to nearby edge nodes (ENs), MEC effectively empowers computation capability and reduces the delay and energy consumption. This improvement enhances the users’ quality of experience (QoE), especially for time-sensitive computation tasks [3], [4].
Efficient task offloading in MEC is a complex optimization challenge due to the dynamic nature of the network and the variety of MDs and servers involved [5], [6]. In particular, determining the optimal offloading strategy, scheduling the tasks, and selecting the most suitable EN for task offloading are the main challenges that demand careful consideration. In addition to the dynamic network changes, the uncertain requirements and sensitive latency properties of computation tasks pose nontrivial challenges that can significantly impact the MEC systems.
To cope with the dynamic nature of the network, recent research has proposed several task offloading algorithms using machine learning methods. In particular, reinforcement learning (RL) [7] holds promises to determine optimal decision-making policies by capturing the dynamics of environments and learning strategies for accomplishing long-term objectives. However, the traditional RL methods are not efficient in handling environments with high-dimensional state spaces. Deep reinforcement learning (DRL) combines traditional RL with deep neural networks to intelligently respond to the unknown and dynamic system when addressing the limitations of RL-based algorithms. Despite these advancements, task offloading still faces significant challenges in real-world scenarios with multiple MDs and ENs. In such scenarios, it is essential for MDs to make offloading decisions independently without prior knowledge of other MDs’ tasks and offloading models. However, existing works fail to adequately address this challenge. In addition, QoE is a time-varying performance measure that reflects user satisfaction and is not affected only by delay, but also by energy consumption. Albeit some existing works have investigated the trade-off between delay and energy consumption, they fail to properly estimate the QoE value and address the users’ requirements. Although DRL offers a valuable tool, a more comprehensive approach is required to accurately estimate the QoE while addressing the aforementioned challenges in practical scenarios where global information is limited.
In this study, we delve into the computation task offloading problem in MEC systems, where strict task processing deadlines and energy constraints can adversely affect the system performance. We propose a distributed QoE-oriented computation offloading (QECO) algorithm to efficiently handle task offloading under uncertain loads at ENs. This algorithm empowers MDs to make offloading decisions utilizing only locally observed information, such as task size, queue details, battery status, and historical workloads at the ENs. To capture the dynamic nature of the MEC environment, we employ the dueling double deep Q-network (D3QN), which is a refined improvement over standard deep Q-network (DQN) model. By integrating both double Q-learning [8] and dueling network architectures [9], D3QN reduces overestimation bias in action-value predictions and more accurately distinguishes the relative importance of states and actions. We also incorporate long short-term memory (LSTM) [10] into our D3QN model to continuously estimate dynamic workloads at ENs. This enables MDs to effectively handle the uncertainty of the MEC environment, where global information is limited, and proactively adjust the offloading strategies to improve their long-term QoE estimation. By adopting the appropriate policy based on each MD’s specific requirements, the QECO algorithm significantly improves the QoE for individual users.
Our main contributions are summarized as follows:
  • Task Offloading Problem in the MEC System: We consider queuing systems for MDs and ENs and formulate the task offloading problem while accounting for the time-varying system environments (e.g., the arrival of new tasks, and the computational requirement of each task). We aim to maximize the long-term QoE, which reflects the task completion rate, task delay, and energy consumption of MDs. Our problem formulation effectively utilizes the resources and properly handles the dynamic fluctuations of workload at ENs.
  • DRL-based Offloading Algorithm: To solve the problem of long-term QoE maximization in highly dynamic mobile environments, we propose the QECO algorithm based on DRL, which empowers each MD to make offloading decisions independently, without prior knowledge of other MDs’ tasks and offloading models. Focusing on the MD’s QoE preference, our approach leverages D3QN and LSTM to prioritize and strike an appropriate balance between QoE factors while accounting for workload uncertainty at the ENs. QECO empowers MDs to anticipate the EN’s load level over time, leading to more accurate offloading strategies. We also analyze the training convergence and computational complexity of the proposed algorithm.
  • Performance Evaluation: We conduct comprehensive experiments to evaluate QECO’s performance and its training convergence against baseline schemes as well as two state-of-the-art existing works. The results demonstrate that our algorithm effectively utilizes the computing resources while taking the dynamic workloads at ENs into consideration. In particular, QECO converges more quickly than vanilla DQN and double DQN (DDQN) methods. QECO also improves average QoE by 37.1% and 29.8% compared to the potential game-based offloading algorithm (PGOA) [11] and distributed and collective DRL-based offloading algorithm (DCDRL) [12], respectively.
The structure of this paper is as follows. Section II reviews the related works. Section III presents the system model, followed by the problem formulation in Section IV. In Section V, we present the algorithm, while Section VI provides an evaluation of its performance. Finally, we conclude in Section VII.
II Related Work
To effectively tackle the challenges of MEC arising from the ever-changing nature of networks, recent research highlights the effectiveness of RL in adapting to environmental changes and learning optimal strategies. In this section, we explore the current RL-based state-of-the-art works and discuss their strengths and limitations. Table I provides an intuitive comparison of these works.
Tabelle I: Comparison of Related Works
Paper Problem Optimization target Method Limitation
Delay Energy Drop QoE
[13] Computation offloading Offline RL Overlooking energy consumption; Not suitable for energy-constrained MEC; May increase energy consumption or reduce the overall system capacity.
[14] Joint offloading and service caching Hierarchical DRL
[15] Computation offloading DQN, DDPG
[16] Computation offloading DDQN
[17] Joint offloading and resource allocation AC network
[18] Computation offloading AC network
[19] Computation offloading Multi-agent RL Do not account for time-sensetive applications; Reducing energy consumption can increase computation delay, potentially resulting in task failures.
[20] Joint offloading and resource allocation DDQN
[21] Joint offloading and resource allocation DDPG
[22] Computation offloading DRL
[23] Privacy aware computation offloading AC network
[24] Computation offloading DDPG Not suitable for applications with strict delay requirements. Tasks may be dropped if there is a strict requirement on either objective.
[25] Online computation offloading DDQN
[26] Computation offloading Multi-agent PPO
[27] Joint offloading and resource allocation DRL
[28] Joint offloading and resource allocation Multi-agent AC
[29] Computation offloading Multi-agent DDPG
[30] Computation offloading Multi-agent AC Overlooking personalized QoE; Treating all MDs similarly when optimzing the overall system performance fails to account for individual user satisfaction.
[31] Privacy preserving offloading Multi-agent AC
[32] Energy harvesting offloading DQN
[33] Joint offloading and resource allocation Multi-agent RL
QECO QoE-oriented computation offloading D3QN + LSTM
For time-sensitive applications, development of delay-optimal approaches is the primary goal. To solve the delay minimization problem in MEC environment, Zhang et al. in [13] proposed an offline RL-based computation offloading algorithm, where the problem is modeled as a repeated game between two agents. Sun et al. in [14] explored both computation offloading and service caching problems in MEC. They formulated an optimization problem that aims to minimize the long-term average service delay. They then proposed a hierarchical DRL framework, which effectively handles both problems under heterogeneous resources. Liu et al. in [15] formulated the offloading problem as an MDP and proposed a deep deterministic policy gradient (DDPG)-based approach to minimize both communication and computation delay in MEC. To minimize total delay and reduce mobile vehicle task waiting time, Tang et al. in [16] developed a dynamic offloading model for multiple vehicles, segmenting tasks into sequential subtasks for more precise offloading decisions. Li et al. in [17] proposed an actor-critic (AC)-based algorithm to optimize computation offloading and resource allocation decisions, which leverages past experience and model knowledge to enable fast and resilient real-time offloading control. In [18], the authors investigated deadline-constrained time-sensitive tasks in vehicular MEC networks and proposed a multi-agent RL-based computation offloading algorithm to address joint delay and computation rate optimization. Despite their benefits, the delay-optimal approaches [13][18] often sacrifice other key performance metrics and may increase energy consumption or diminish overall computational capacity.
For energy-constrained MEC systems, reducing energy consumption becomes the key objective, motivating the development of energy-efficient approaches. Munir et al. in [19] investigated an energy dispatch problem for a self-powered MEC system and proposed a semi-distributed approach using a multi-agent RL framework to minimize energy consumption. Zhou et al. in [20] proposed a Q-learning approach, which is an extension of RL to achieve optimal resource allocation strategies and computation offloading. Dai et al. in [21] introduced the integration of action refinement into DRL and designed an algorithm based on DDPG to optimize resource allocation and computation offloading concurrently. To ensure that minimizing energy consumption does not lead to a reduction in the system’s computing rate, researchers have expanded their work by incorporating maximum tolerable delay constraints. In [22], Chouikhi et al. proposed a DRL-based computation offloading approach that aims to minimize long-term energy consumption and maximize the number of tasks completed within their deadlines. To address the privacy-aware computation offloading problem, Wu et al. in [23] proposed a DQN-based method to optimize the computation rate and energy consumption in a queuing-based Industrial Internet of Things (IIoT) network. Huang et al. in [24] proposed a DRL-based method using partially observable MDP (POMDP), which ensures the real-time tasks deadlines are met while minimizing the total energy consumption of MDs. This algorithm effectively tackles the challenges of dynamic resource allocation in large-scale heterogeneous networks. The aforementioned energy-efficient approaches [19][24] do not adequately address the requirements of time-sensitive applications, where reducing delay could have significant effects on the users’ QoE.
Energy consumption and computation delay are two conflicting objectives that must be properly balanced when dealing with time-sensitive applications in energy-constrained MEC systems. Liao et al. in [25] introduced a DDQN-based algorithm for performing online computation offloading in MEC. This algorithm optimizes transmission power and CPU frequency when minimizing both task computation delay and energy consumption. To optimize delay and energy consumption, Wu et al. in [26] investigated the computation offloading problem in a queuing-based MEC IIoT system. They modeled the problem as a POMDP and proposed a multi-agent proximal policy optimization (PPO)-based method to obtain the optimal offloading strategy in dynamic environments. Gong et al. in [27] proposed a DRL-based algorithm jointly optimizes task offloading and resource allocation to achieve lower energy consumption and decreased task delay in IIoT systems. Liu et al. in [28] investigated a two-timescale computing offloading and resource allocation problem and proposed a resource coordination algorithm based on multi-agent DRL, which can generate interactive information along with resource allocation decisions. She et al. in [29] formulated the computation offloading problem as a POMDP and proposed a DRL-based approach to minimize both transmission delay and energy. Although delay-energy trade-off approaches [25][29] effectively address the needs of time-sensitive tasks, they may lead to task failures and drop when strict requirements are imposed on one of the objectives.
In addition to energy and delay, task computation rate is also of paramount importance. An efficient offloading approach should jointly consider these factors and achieve a proper trade-off among them. In [30], Gao et al. introduced an attention-based multi-agent algorithm designed for decentralized computation offloading that jointly considers the aforementioned objectives. To optimize privacy protection and quality of service, authors in [31] investigated the joint computation offloading and power allocation problems for an IIoT network. They modeled the problem as an MDP and proposed a multi-agent DQN-based algorithm. In [32], the authors proposed an offloading algorithm using DQN for wireless-powered IoT devices in MEC. This algorithm aims to minimize the task drop rate while the devices solely rely on harvested energy for operation. Wu et al. in [33] introduced a stochastic game-based resource allocation in the MEC. They used an MDP and proposed a multi-agent RL method with the goal of minimizing both energy consumption and delay. Studies [30][33] have examined delay-sensitive applications with processing deadlines and addressed energy-constrained MEC challenges, but they overlook user-specific requirements and fail to effectively meet QoE expectations.
The aforementioned existing works have advanced computation offloading in MEC systems by targeting various objectives, which are usually conflicting. These approaches typically compromise one or a subset of objectives to meet their targets. However, while promising, they may not be suitable for time-sensitive applications in energy-constrained MEC environments, especially under the stringent requirements of MDs and the highly dynamic loads of ENs. For instance, studies [13][24] focus solely on energy consumption or computation delay. While [25][29] jointly consider these two targets, they neglect task completion and fail to prevent task failures. Recent works [30][33] attempt to address these challenges by incorporating energy, delay, and task completion. However, they disregard users’ QoE requirements and their individual preferences. QoE is a time-varying performance measure that reflects user satisfaction, with each user prioritizing its specific needs over overall system performance. These gaps motivate us to design an adaptive algorithm to balance conflicting objectives and enhance QoE for MEC users individually.
In addition to the above limitations, the constrained computing recourses of ENs necessitate careful attention. Existing works either neglect the resource constraints (e.g., [34]) or fail to properly account for the ENs’ dynamic load, particularly under limited information (e.g., [14], [21], [27], [30], [31], [35]). These lead to degraded system performance. Different from these works, we propose a distributed algorithm that effectively manages unknown load dynamics at ENs. Our algorithm enables each MD to make offloading decisions independently, without requiring information (e.g., task models, offloading decisions) from other MDs, thereby reducing signaling overhead and improving overall system performance.
III System Model
We investigate a MEC system consisting of a set of MDs denoted by $\mathcal{I}=\{1,2,...,I\}$, along with a set of ENs denoted by $\mathcal{J}=\{1,2,...,J\}$, where $I$ and $J$ represent the number of MDs and ENs, respectively. We regard time as a specific episode containing a series of $T$ time slots denoted by $\mathcal{T}=\{1,2,\ldots,T\}$, each representing a duration of $\tau$ seconds. As shown in Fig. 1, we consider two separate queues for each MD to organize tasks for local processing or dispatching to ENs, operating in a first-in-first-out (FIFO) manner. The MD’s scheduler is responsible for assigning newly arrived tasks to each of the queues at the beginning of the time slot. On the other hand, we assume that each EN $j\in\mathcal{J}$ consists of $I$ FIFO queues, where each queue corresponds to an MD $i\in\mathcal{I}$. When each task arrives at an EN, it is enqueued in the corresponding MD’s queue.
We define $z_{i}(t)$ as the index assigned to the computation task arriving at MD $i\in\mathcal{I}$ in time slot $t\in\mathcal{T}$. Let $\lambda_{i}(t)$ denote the size of this task in bits. The size of task $z_{i}(t)$ is selected randomly and uniformly from a discrete set $\Lambda=\{\lambda_{1},\lambda_{2},\ldots,\lambda_{\theta}\}$, where $\theta$ represents the number of these values. Note that task sizes are drawn from a discrete set since, in many applications, tasks typically come with predefined sizes, [36], [37]. We consider $\lambda_{i}(t)\in\Lambda\cup\{0\}$ to include the case that no task has arrived. We also denote the task’s processing density as $\rho_{i}(t)$ that indicates the number of CPU cycles required to complete the execution of a unit of the task. Furthermore, we denote the deadline of this task by $\Delta_{i}(t)$ which is the number of time slots that the task must be completed to avoid being dropped.
III-A Communication Model
$\delta_{i}^{\text{T}}(t)={\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}i}\left[{\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}i}\max\limits_{t^{\prime}{\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}i}\in{\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}i}\{0,1,\ldots,t-1\}}l_{i}^{\text{T}}{\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}i}(t^{\prime})-t+1{\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}i}\right]^{+}{\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}i},$ (1)
@@ III-A Communication Model
$D_{i}^{\text{T}}(t)=\sum_{j\in\mathcal{J}}y_{i,j}(t){\lambda_{i}(t)\over r_{i,j}(t)\tau}.$ (3)